OpenStudy (anonymous):

Use Newton's method to approximate a root of the equation cos(x^2 - 7) = x^3. Let x1 = 1 be the initial approximation.

OpenStudy (campbell_st):

the equation needs to be rewritten \[f(x) = \cos(x^2 - 7) - x^3 \] the derivative is \[f'(x) = -2x \sin(x^2 - 7) - 3x^2\] now use the formula \[x _{n + 1} = x _{n} - f(x _{n})/f'(x _{n})\] and \[x _{1} = 1\]